# Deep Learning based Visually Rich Document Content Understanding: A Survey

### v1

@@ 5.2. Visual Question Answering
Name Conf./J. Year Domain Lang. # Docs # Images # Questions Answer Type MP Format Metrics Annotation
DocVQA WACV 2021 Industrial Reports English N/A 12,767 50,000 Text N D./P./H. ANLS Human
VisualMRC AAAI 2021 Website English N/A 10,197 30,562 Text N D. BLUE, etc Human
TAT-DQA MM 2022 Financial Reports English 2,758 3,067 16,558 Text/RS-Gen. Y D. EM, F1 Human
RDVQA MM 2022 Data Analysis Report N/A 8,362 8,514 41,378 Text N D. ANLS, ACC Human
CS-DVQA MM 2022 Industry Documents English N/A 600 1,000 Text and Nodes N D./P./H. ANLS Human
PDFVQA-Task A ECML-PKDD 2023 Academic Paper English N/A 12,337 81,085 Num or Yes/No N D. F1 Template
PDFVQA-Task B ECML-PKDD 2023 Academic Paper English N/A 12,337 53,872 Entity N D. F1 Template
PDFVQA-Task C ECML-PKDD 2023 Academic Paper English 1,147 12,337 5,653 Entity Y D. EM Template
MPDocVQA PR 2023 Industrial Reports English 6,000 48,000 46,000 Text Y D./P./H. ANLS Human
DUDE ICCV 2023 Cross-domain English 5,019 28,709 41,541 Text, Yes/No Y D. ANLS Human
MMVQA IJCAI 2024 Academic Paper English 3,146 30,239 262,928 Entity Y D. EM, PM, MR LLM + Human
Table 2. Summary of visually rich document question answering datasets
5.2.1. Single Page VRD-QA Datasets
@@
DocVQA (Mathew et al., 2021) is a pioneering dataset in document-based Visual Question Answering (VQA), sourced from the UCSF Industry Document Library. It comprises 50,000 manually generated questions framed on 12,767 document images, encompassing digital, printed, and handwritten formats. The dataset follows an extractive-style QA format similar to benchmarks like SQuAD (Rajpurkar et al., 2016b) and VQA (Biten et al., 2019b). Evaluation typically involves fine-grained models such as LayoutLM variants (Xu et al., 2020a, b; Huang et al., 2022), LiLT (Wang et al., 2022b), and generative models (Tang et al., 2022; Kim et al., 2022), using metrics like Average Normalized Levenshtein Similarity (ANLS) (Biten et al., 2019a). However, it requires additional processing for coarse-grained models like SelfDoc (Li et al., 2021b). CS-DVQA (Du et al., 2022) builds upon DocVQA by enhancing QA pairs to better reflect real-world requirements. It extracts 600 images from the DocVQA dataset and generates 1,000 QA pairs under human supervision. During question generation, it incorporates common-sense knowledge from real life, expanding answers beyond extractive in-line text to include question-related nodes (Nodes) sourced from ConceptNet (Speer et al., 2017).
@@
VisualMRC (Tanaka et al., 2021) is compiled from website screenshots across 35 domains, carefully selected to exclude pages with handwritten content and to prefer pages containing short text (no more than 2 to 3 paragraphs). Unlike other datasets that might only provide question-answer annotations (Mathew et al., 2021) or automatically acquire document semantic entities (Ding et al., 2023b), VisualMRC includes manually annotated layout structures with fine-grained semantic entity types such as Heading, Paragraph, Subtitle, Picture, and Caption. Question-answer pairs are generated through crowdsourcing. Consequently, VisualMRC is well-suited for evaluating both fine-grained and coarse-grained-based QA frameworks, providing a rich resource for assessing the effectiveness of models in understanding and interpreting detailed document layouts and semantic entities.
@@
PDFVQA-Task A and Task B (Ding et al., 2023b) form part of the first document VQA dataset from PubMed Central, focusing on content and structural understanding. This dataset includes three tasks: two for single-page documents (Tasks A and B) and one for multi-page documents (Task C). Task A evaluates the structural and spatial relationships within document images, with answers being either counts or Yes/No. Task B focuses on extracting document entities based on their logical and spatial configurations. The PDFVQA dataset provides only coarse-grained, entity-level annotations, necessitating further processing for models that require fine-grained analysis. This setup is ideal for testing models’ capabilities in understanding the logical and spatial structures of document images.
5.2.2. Multi-Page VRD-QA Datasets
@@
TAT-DQA (Zhu et al., 2022), an extension of the TAT-QA (Zhu et al., 2021a) dataset, is developed with more complex natural document structures and an expanded set of manually corrected and generated question-answer pairs derived from business financial reports. Unlike other datasets that primarily focus on extractive or simple abstractive answers (such as counting or yes/no), TAT-DQA includes questions requiring arithmetic reasoning, where values must be extracted from tabular data and textual content for discrete calculations. This dataset adopts the evaluation metrics of TAT-QA, including Exact Matching and a numeracy-focused F1 score. These metrics are particularly tailored to assess the accuracy of arithmetic reasoning and data extraction capabilities of the models tested with TAT-DQA.
@@
RDVQA dataset (Wu et al., 2022) compiles a large collection of conversational chats and associated images from an E-commerce platform. It employs standard OCR and Named Entity Recognition (NER) techniques to extract text and redact sensitive information, ensuring privacy protection through masking. The dataset includes question-answer pairs within the images, which are manually verified to confirm image clarity and the presence of at least one question-answer pair per image. Although some documents span multiple pages, this dataset is structured such that it can be processed relatively easily by single-page VRD-QA models.
@@
PDFVQA-Task C (Ding et al., 2023b) is a distinct sub-task within the PDFVQA dataset that expands document VQA to encompass entire long documents, moving beyond the single-page focus of Tasks A and B. In Task C, to answer questions, the model often needs to retrieve information from multiple document entities. Thus, Task C employs Exact Matching for its ground truth annotations. Similar to Tasks A and B, additional processing is required for evaluating models at a fine-grained level.
@@
MP-DocVQA (Tito et al., 2023) extends the original DocVQA (Mathew et al., 2021) dataset to accommodate multi-page document analysis. This version includes adjacent pages from the same documents, expanding the dataset from 12,767 to 64,057 document images. In adapting to a multi-page format, some questions inappropriate were removed. However, it’s important to note that while the dataset allows for questions across multiple pages, the answers remain confined to individual pages; there are no cross-page answers in the MP-DocVQA dataset.
@@
DUDE (Van Landeghem et al., 2023) is the first cross-domain, multi-page document VQA dataset, featuring a diverse collection of documents from various fields such as medical, legal, technical, and financial, and different document types including CVs, reports, and papers. It comprises 5,019 documents, 28,709 document pages, and 41,541 manually annotated questions. Question types vary from extractive in-line text and Yes/No answers to multi-hop reasoning and structural understanding, similar to those in PDFVQA (Ding et al., 2023b). To evaluate model performance, DUDE uses the ANLS metric (Mathew et al., 2021) for assessing answer prediction accuracy. Additionally, it employs two other metrics: Expected Calibration Error (Guo et al., 2017) and Area-Under-Risk-Coverage-Curve (CURC) (Geifman and El-Yaniv, 2017; Jaeger et al., 2022) to gauge the overconfidence and miscalibration in document understanding models. These features make DUDE a comprehensive tool for evaluating cross-domain document understanding models.
@@
MMVQA (Ding et al., 2024a) is a dataset sourced from PubMed Central, designed for the retrieval of multimodal semantic entities from multi-page documents. The questions are generated using ChatGPT (OpenAI, 2023) and subsequently verified manually. Unlike other datasets that focus solely on in-line text or text-dense entities, MMVQA also considers entire tables and figures as potential answers to the given questions. This dataset introduces various evaluation metrics to cater to different application scenarios: Exact Matching and Partial Matching Accuracy assess the precision of responses, while Multi-label Recall evaluates how well the model identifies all relevant answers across the document. This diverse set of metrics makes MMVQA suitable for comprehensive performance evaluation in complex, multimodal document understanding tasks.
5.3. Summary of VRDU Datasets
The common characteristics of KIE and VRD-QA datasets are represented in Table 1 and Table 2. Most existing benchmark datasets for key information extraction are designed for single-page scenarios and predominantly cater to English documents. However, real-world applications often involve more complex multi-page forms. Even forms typically require input from multiple parties and contain multiple languages, such as customs or import/export declaration forms, presenting unique challenges not addressed by current datasets. Similarly, in the domain of VRD-QA, while multi-page VRD-QA has gained increased attention, the integration and exploration of multimodal information remain insufficiently developed. Specifically, structure and relation-aware VRD-QA, which is crucial for interpreting and understanding cross-page relationships, is still a largely unexplored area.
6. Critical Discussion
In Section 3 and Section 3, the models proposed for handling both mono-task and multi-task Visually Rich Document Understanding (VRDU) are introduced. This section aims to provide a summary and analysis of relevant techniques, including their application scenarios, advantages, and disadvantages. Several key topics are covered: Feature Engineering, Cross-Modality Fusion, Model Architecture, and Pre-training Mechanisms. Additionally, with the rapid development of Large Language Models (LLMs), LLM-based frameworks and emerging trends in applying LLMs to VRDU are discussed.
6.1. Feature Representation
6.1.1. Textual Representation
@@
Text in VRDs provides essential semantic context, crucial for understanding content and conducting various downstream tasks. Depending on the information granularity required by the framework and application scenarios, textual representation methods can generally be categorized at the word or entity level.
Figure 4. Fine-grained (word-level) and coarse-grained (entity-level) textual information encoding for VRDU frameworks.
@@
Word-level Representations In Visually Rich Documents (VRDs), text sequences extracted by off-the-shelf OCR tools or PDF parsers (e.g., PDFMiner) can be encoded using word embedding methods such as Word2Vec (Mikolov et al., 2013a) and Glove (Pennington et al., 2014). For more comprehensive textual embeddings, various BERT-style bi-directional pretrained transformer models like BERT (Devlin, 2018) and RoBERTa (Liu et al., 2019b) are employed to generate context-aware word representations. As the visually rich and structurally complex nature of VRDs, layout-aware and visual-integrated fine-grained models have been developed, such as LayoutLM families (Xu et al., 2020a, b; Huang et al., 2022; Xu et al., 2021), LiLT (Wang et al., 2022b). These models generate word representations that integrate multimodal information, combining text, visual cues, and layout structure to achieve SoTA performance on several downstream tasks.
@@
Entity-level Representations To acquire a dense representation of a text sequence for performing entity-level VRDU tasks, various approaches are adopted. These include averaging word embeddings specific to an entity or leveraging the $<CLS>$ token to encapsulate the entire sequence, including averaging the word embeddings belonging to an entity or using $<CLS>$ token to represent an entire sequence. SentenceBERT (Reimers and Gurevych, 2019) is also often adopted to encode text sequences within document entities. However, a standardized approach for acquiring textual representations of entities is yet to be established, necessitating preliminary testing and validation.
6.1.2. Visual Representation
Figure 5. Commonly adopted visual information encoding approaches.

### v2

@@ 5.2 Visually Rich Document Question Answering
Name Venue Year Domain # Doc. # Img # Q Answer Type MP Format Metrics Anno.
DocVQA WACV 2021 Industrial Reports N/A 12,767 50,000 Text N D./P./H. ANLS Human
VisualMRC AAAI 2021 Website N/A 10,197 30,562 Text N D. BLUE, etc Human
TAT-DQA MM 2022 Financial Reports 2,758 3,067 16,558 Text/RS-Gen. Y D. EM, F1 Human
RDVQA MM 2022 Data Analysis Report 8,362 8,514 41,378 Text N D. ANLS, ACC Human
CS-DVQA MM 2022 Industry Documents N/A 600 1,000 Text and Nodes N D./P./H. ANLS Human
PDFVQA-Task A PKDD 2023 Academic Paper N/A 12,337 81,085 Num or Yes/No N D. F1 Template
PDFVQA-Task B PKDD 2023 Academic Paper N/A 12,337 53,872 Entity N D. F1 Template
PDFVQA-Task C PKDD 2023 Academic Paper 1,147 12,337 5,653 Entity Y D. EM Template
MPDocVQA PR 2023 Industrial Reports 6,000 48,000 46,000 Text Y D./P./H. ANLS Human
DUDE ICCV 2023 Cross-domain 5,019 28,709 41,541 Text, Yes/No Y D. ANLS Human
MMVQA IJCAI 2024 Academic Paper 3,146 30,239 262,928 Entity Y D. EM, PM, MR LLM + H
Table 2: Summary of visually rich document question answering (VRD-QA) datasets. MP: Multi-page; D./P./H.: Digital / Printed / Handwritten formats; RS-Gen.: Reasoning-based Generation; Metrics: ANLS = Average Normalized Levenshtein Similarity, EM = Exact Match, F1 = F1-score, PM = Partial Match, MR = Multilabel Recall, ACC = Accuracy, BLEU = BLEU score; Anno.: Annotation source (Human, Template, or LLM + Human).
5.2.1 Single Page VRD-QA Datasets
@@
DocVQA [docvqa] is a pioneering dataset in document-based Visual Question Answering (VQA), sourced from the UCSF Industry Document Library. It comprises 50,000 manually generated questions framed on 12,767 document images, encompassing digital, printed, and handwritten formats. The dataset follows an extractive-style QA format similar to benchmarks like SQuAD [rajpurkar2016squad] and VQA [biten2019scene]. Evaluation typically involves fine-grained models such as LayoutLM variants [layoutlm, layoutlmv2, layoutlmv3], LiLT [lilt], and generative models [udop, donut], using metrics like Average Normalized Levenshtein Similarity (ANLS) [stvqa]. However, it requires additional processing for coarse-grained models like SelfDoc [selfdoc]. CS-DVQA [calm] builds upon DocVQA by enhancing QA pairs to better reflect real-world requirements. It extracts 600 images from the DocVQA dataset and generates 1,000 QA pairs under human supervision. During question generation, it incorporates common-sense knowledge from real life, expanding answers beyond extractive in-line text to include question-related nodes (Nodes) sourced from ConceptNet [speer2017conceptnet].
@@
VisualMRC [visualmrc] is compiled from website screenshots across 35 domains, carefully selected to exclude pages with handwritten content and to prefer pages containing short text (no more than 2 to 3 paragraphs). Unlike other datasets that might only provide question-answer annotations [docvqa] or automatically acquire document semantic entities [pdfvqa], VisualMRC includes manually annotated layout structures with fine-grained semantic entity types such as Heading, Paragraph, Subtitle, Picture, and Caption. Question-answer pairs are generated through crowdsourcing. Consequently, VisualMRC is well-suited for evaluating both fine-grained and coarse-grained-based QA frameworks, providing a rich resource for assessing the effectiveness of models in understanding and interpreting detailed document layouts and semantic entities.
@@
PDFVQA-Task A and Task B [pdfvqa] form part of the first document VQA dataset from PubMed Central, focusing on content and structural understanding. This dataset includes three tasks: two for single-page documents (Tasks A and B) and one for multi-page documents (Task C). Task A evaluates the structural and spatial relationships within document images, with answers being either counts or Yes/No. Task B focuses on extracting document entities based on their logical and spatial configurations. The PDFVQA dataset provides only coarse-grained, entity-level annotations, necessitating further processing for models that require fine-grained analysis. This setup is ideal for testing models’ capabilities in understanding the logical and spatial structures of document images.
5.2.2 Multi-Page VRD-QA Datasets
@@
TAT-DQA [tatdqa], an extension of the TAT-QA [zhu2021tat] dataset, is developed with more complex natural document structures and an expanded set of manually corrected and generated question-answer pairs derived from business financial reports. Unlike other datasets that primarily focus on extractive or simple abstractive answers (such as counting or yes/no), TAT-DQA includes questions requiring arithmetic reasoning, where values must be extracted from tabular data and textual content for discrete calculations. This dataset adopts the evaluation metrics of TAT-QA, including Exact Matching and a numeracy-focused F1 score. These metrics are particularly tailored to assess the accuracy of arithmetic reasoning and data extraction capabilities of the models tested with TAT-DQA.
@@
RDVQA dataset [rdvqa] compiles a large collection of conversational chats and associated images from an E-commerce platform. It employs standard OCR and Named Entity Recognition (NER) techniques to extract text and redact sensitive information, ensuring privacy protection through masking. The dataset includes question-answer pairs within the images, which are manually verified to confirm image clarity and the presence of at least one question-answer pair per image. Although some documents span multiple pages, this dataset is structured such that it can be processed relatively easily by single-page VRD-QA models.
@@
PDFVQA-Task C [pdfvqa] is a distinct sub-task within the PDFVQA dataset that expands document VQA to encompass entire long documents, moving beyond the single-page focus of Tasks A and B. In Task C, to answer questions, the model often needs to retrieve information from multiple document entities. Thus, Task C employs Exact Matching for its ground truth annotations. Similar to Tasks A and B, additional processing is required for evaluating models at a fine-grained level.
@@
MP-DocVQA [mpdocvqa] extends the original DocVQA [docvqa] dataset to accommodate multi-page document analysis. This version includes adjacent pages from the same documents, expanding the dataset from 12,767 to 64,057 document images. In adapting to a multi-page format, some questions inappropriate were removed. However, it’s important to note that while the dataset allows for questions across multiple pages, the answers remain confined to individual pages; there are no cross-page answers in the MP-DocVQA dataset.
@@
DUDE [dude] is the first cross-domain, multi-page document VQA dataset, featuring a diverse collection of documents from various fields such as medical, legal, technical, and financial, and different document types including CVs, reports, and papers. It comprises 5,019 documents, 28,709 document pages, and 41,541 manually annotated questions. Question types vary from extractive in-line text and Yes/No answers to multi-hop reasoning and structural understanding, similar to those in PDFVQA [pdfvqa]. To evaluate model performance, DUDE uses the ANLS metric [docvqa] for assessing answer prediction accuracy. Additionally, it employs two other metrics: Expected Calibration Error [guo2017calibration] and Area-Under-Risk-Coverage-Curve (CURC) [geifman2017selective, jaeger2022call] to gauge the overconfidence and miscalibration in document understanding models. These features make DUDE a comprehensive tool for evaluating cross-domain document understanding models.
@@
MMVQA [mmvqa] is a dataset sourced from PubMed Central, designed for the retrieval of multimodal semantic entities from multi-page documents. The questions are generated using ChatGPT [chatgpt] and subsequently verified manually. Unlike other datasets that focus solely on in-line text or text-dense entities, MMVQA also considers entire tables and figures as potential answers to the given questions. This dataset introduces various evaluation metrics to cater to different application scenarios: Exact Matching and Partial Matching Accuracy assess the precision of responses, while Multi-label Recall evaluates how well the model identifies all relevant answers across the document. This diverse set of metrics makes MMVQA suitable for comprehensive performance evaluation in complex, multimodal document understanding tasks.
5.2.3 Summary of VRD-QA Datasets
Similar to KIE datasets, in Table 2, the development of VRD-QA datasets also demonstrates a marked progression toward practical and challenging scenarios. While earlier datasets [docvqa, visualmrc] concentrated on single-page, clean digital documents with simple textual answers, newer benchmarks increasingly demand complex reasoning over multi-page layouts and multimodal semantic entities (e.g., MPDocVQA [mpdocvqa], MMVQA [mmvqa]), accommodate diverse and domain-specific content (e.g., financial, academic, cross-domain), and support mixed formats including printed and handwritten inputs. The variety in answer formats—ranging from span-based text to entities, binary decisions, and reasoning-based generation—combined with rich evaluation protocols (e.g., EM, F1, ANLS, BLEU), underscores the rising complexity and real-world applicability of the task, pushing models toward deeper semantic comprehension, robust layout interpretation, and multimodal reasoning under noisy and heterogeneous conditions.
5.3 Evaluation Metrics
F1-score The F1-score is a widely used evaluation metric in KIE, VQA, entity recognition, and document classification. It helps assess how well a model can identify relevant elements while minimizing false positives and false negatives. It provides a balanced measure of a model’s precision and recall, calculated as the harmonic mean of precision and recall using the following formula:
F1 score = 2 × (Precision × Recall) / (Precision + Recall)
5.3.1 Evaluation Metrics in KIE Datasets
AP (Average Precision) AP measures the area under the precision–recall (P–R) curve for ranked retrieval results, calculating the average of precision values at different recall levels. In document understanding and object detection tasks, AP measures how accurately a model can identify and localize different document elements such as tables, figures, and text blocks.
CLEval (Character-Level Evaluation) CLEval [baek2020cleval] is designed for text detection and recognition tasks, providing a fine-grained assessment through: instance matching process and character scoring process. The instance matching process handles granularity differences between predicted and ground-truth text regions, and the character-level scoring process measures partial correctness. CLEval evaluation metrics provides more nuanced evaluation by considering partially correct results.
5.3.2 Evaluation Metrics in VRD-QA Datasets
ANLS (Average Normalized Levenshtein Similarity) ANLS [biten2019icdar] is a metric designed to evaluate the similarity between predicted and ground truth answers, particularly in tasks such as VQA, where outputs may contain minor errors such as typos. Formally, ANLS can be defined as: For $N$ questions, each with $M$ ground truth answers $a_{ij}$ and a prediction $o_{q_{i}}$
$\mathrm{ANLS}=\frac{1}{N}\sum_{i=1}^{N}\left(\max_{j}s(a_{ij},o_{q_{i}})\right)$ (14)
where $s(a_{ij},o_{q_{i}})$ is the thresholded normalized similarity for the $j$-th ground-truth to the $i$-th prediction. ANLS normalizes Levenshtein distance to account for differences in string lengths, producing a score between 0 and 1, where 1 indicates perfect similarity. ANLS gives a high score when predicted text is very similar to the ground truth and gracefully degrades as errors increase.
BLEU (Bilingual Evaluation Understudy) The BLEU score is a metric originally designed for evaluating machine translation quality, but has been adapted for various text generation tasks, including VQA. BLEU measures the similarity between machine-generated text and reference texts by comparing n-grams. BLEU scores range from 0 to 1, with higher scores indicating greater similarity to reference texts.
ACC (Accuracy) Accuracy measures the proportion of correctly predicted instances over the total number of instances. In VRD-QA tasks [docvqa], this typically refers to the percentage of questions where the predicted answer exactly matches at least one ground-truth answer. A stricter variant, Exact Match (EM), scores an instance as correct only if the predicted entity set is identical to the ground-truth set [mmvqa]. Conversely, Partial Match (PM) is a more lenient metric that gives credit for partial correctness—an instance is considered correct if there is any non-empty overlap between the predicted and ground-truth entity sets. While EM emphasizes exact retrieval, PM is more suitable for scenarios where partial information suffices, such as identifying at least one key entity within a document.
MR (Multilabel Recall) MR is a metric designed for multilabel classification tasks where each instance can belong to multiple categories simultaneously. This metric evaluates the proportion of actual positive labels correctly identified by the model and is particularly important in situations where identifying all relevant positive instances is critical. MR is calculated as: MR = True Positives / (True Positives + False Negatives)
6 Technique Overview with Performance Analysis
Sections 3 and 3 present models designed for mono-task and multi-task VRD-CU. This section reviews core components—Feature Representation, Cross-Modality Fusion, Model Architecture, and Pre-training—with a focus on their functional roles, advantages, and limitations across different scenarios. It further discusses recent progress in integrating LLM and MLLM into VRD-CU pipelines, highlighting emerging trends and paradigm shifts. Finally, to provide a comprehensive understanding of model performance, we summarize the quantitative results of representative approaches four broadly used VRD-CU benchmarks, offering a comparative view of their empirical effectiveness.
6.1 Feature Representation
6.1.1 Textual Representation
@@
Text in VRDs provides essential semantic context, crucial for understanding content and conducting various downstream tasks. Depending on the information granularity required by the framework and application scenarios, textual representation methods can generally be categorized at the word or entity level, as illustrated by Figure 4.
Figure 4: Fine-grained vs. coarse-grained textual embeddings in VRD-CU. Left: Word-level embedding uses $\mathcal{E}(\text{text})=\{t_{1},t_{2},\dots,t_{n}\}$. Right: Entity-level embedding aggregates token vectors via mean pooling: $e=\frac{1}{k}\sum_{i=1}^{k}t_{i}$, where $\{t_{1},\dots,t_{k}\}$ are tokens within an entity.
@@
Word-level Representations In VRDs, text sequences extracted by off-the-shelf OCR tools or PDF parsers (e.g., PDFMiner) can be encoded using word embedding methods such as Word2Vec [word2vec] and Glove [glove]. For more comprehensive textual embeddings, various BERT-style bi-directional pretrained transformer models like BERT [bert] and RoBERTa [roberta] are employed to generate context-aware word representations. As the visually rich and structurally complex nature of VRDs, layout-aware and visual-integrated fine-grained models have been developed, such as LayoutLM families [layoutlm, layoutlmv2, layoutlmv3, layoutxlm], LiLT [lilt]. These models generate word representations that integrate multimodal information, combining text, visual cues, and layout structure to achieve SoTA performance on several downstream tasks.
@@
Entity-level Representations To acquire a dense representation of a text sequence for performing entity-level VRD-CU tasks, various approaches are adopted. These include averaging word embeddings specific to an entity or leveraging the [CLS] token to encapsulate the entire sequence, including averaging the word embeddings belonging to an entity or using [CLS] token to represent an entire sequence. SentenceBERT [sentencebert] is also often adopted to encode text sequences within document entities. However, a standardized approach for acquiring textual representations of entities is yet to be established, necessitating preliminary testing and validation.
6.1.2 Visual Representation
Figure 5: Comparison of visual backbone architectures for document understanding. Left: CNN-based backbones extract features at the entity/token or patch level using region-of-interest (RoI) pooling. Right: Vision Transformer (ViT)-style encoders operate directly on fixed-size image patches, enabling patch-level feature extraction across the entire document. Both approaches utilize pretrained models to obtain visual embeddings $\{v_{1},v_{2},\ldots\}$ for downstream tasks.